Let U and W be subspaces of a vector space V of finite dimension. Prove that(UnW)°=U°+W°


1
Expert's answer
2021-12-23T08:44:12-0500

To prove this, three things must be ascertained:




i. It must not be empty.




ii. The sum of the two sets must hold.




iii. The scalar multiplication of the two sets must hold.


i) 0=0+0, elements of U + W

ii) ß(U + W)= ẞU + ẞW, elements of U+W

iii) (U1+W1) + (U2+W2)= (U1+U2) +(W1+W2), elements of U+W

https://youtu.be/Rcj1-E3SAhs